# Kolmogorov-Arnold Energy Models: Fast and Interpretable Generative Modeling

### v8

@@ 2.3.1 Generator
Thus, a smooth transition between the latent space and data space can be achieved by progressively doubling layer widths, eliminating the need for arbitrary choices while maintaining relaxed compliance with KART. Strict adherence to KART can be achieved for hidden layer $h_{l_{k},l_{k-1}}$ by applying either a sigmoid or affine transformation after layer $h_{l_{k-2},l_{k-3}}$ to shift its range back to $0,1]^{L_{k-1}}$. However, using sigmoid may lead to vanishing gradients and we suggest that function domains are less critical to KART than its structural properties and should not be prioritized over maintaining gradient flow and avoiding saturation.
Adopting a KAN-based generator also allows for training with L1 regularization, followed by pruning with the method proposed by liu2024kankolmogorovarnoldnetworks. This enables further model compression beyond the baseline structure defined by KART.
2.3.2 Mixture prior
As noted in Sec. 2.1, the independence of the priors $\pi_{q,p}$ limits expressivity. Moreover, the summation over $p$ in Eq. 9 accumulates the variances of sampled $z_{q,p}$, which can lead to overly diffuse representations. To address this, particularly for more complex generative tasks such as those in Sec. 3.2, we relax the strict adherence to KART by reformulating the prior as a mixture distribution, similar to the approach described by rivera2024trainvae. This involves moving the summation over p{p} from 1 to $n_{z}$ out of the generator and absorbing it into the prior, yielding a mixture density:
$p_{\bm{f},q}(z)$ $=\sum_{p=1}^{n_{z}}\alpha_{q,p}\frac{\exp(f_{q,p}(z))}{Z_{q,p}}\pi_{0}(z),\quad\text{where}\quad\sum_{p=1}^{n_{z}}\alpha_{q,p}=1,\quad\alpha_{q,p}\geq 0.$ (13)
Here, $\alpha_{q,p}$ represent the mixture proportions, which must satisfy non-negativity and unit-sum constraints. In this study, they were fixed uniformly as $1/n_{z}$, guided by rivera2024trainvae. They could alternatively be learned by passing trainable parameters through a softmax function, (to enforce the constraints), however this approach was not used for the experiments in Sec. 3.
The result is a mixture of energy-based priors. Sampling becomes more efficient than in the previous formulation, as it proceeds component-wise using ITS for discrete variables, (DEVROYE200683). However, this violates KART, since $F$ now corresponds to a discrete rather than continuous CDF.
The mixture proportions are learned by replacing the log-prior in Eq. 18 with the logarithm of Eq.13. To encourage balanced weighting and reduce model complexity, L1 regularization can be applied to $\alpha_{q,p}$ and small values can be pruned. Component selection follows a categorical distribution over $p=1,\dots,n_{z}$, determined by $\alpha_{q,p}$. Let $u_{\text{mix}}\sim\text{Unif}(0,1)$. Guided by DEVROYE200683 regarding how ITS applied to discrete variables, the selected component $p^{*}$ is given by:
$p^{*}=\min\left\{j\mid\sum_{p=1}^{j}\alpha_{q,p}\geq u_{\text{mix}}\right\}.$ (14)
This ensures that each component $p^{*}$ is selected with probability $\alpha_{q,p^{*}}$. ITS can then be used to sample from the chosen component, (see  2.4.3).
2.4 Maximum likelihood training
With $N_{x}$ training examples in a batch, the optimization objective is the minimization of the negative log-marginal likelihood:
$\min_{\bm{f},\bm{\Phi}}\mathcal{L}(\bm{x}^{(1)},\dots,\bm{x}^{(N_{x})})=\min_{\bm{f},\bm{\Phi}}\left[-\sum_{b=1}^{N_{x}}\log\left(p_{\bm{f},\mathbf{\Phi}}(\bm{x}^{(b)})\right)\right].$ (15)
The dependency on the collection, $\bm{f},\mathbf{\Phi}$, represents all learnable elements of KAEM. For each sample of $\bm{x}^{(b)}$, the log-marginal likelihood in Eq. 15 can be separated out into two log-densities by first introducing the latent vector, $\bm{z}$, otherwise absent upon marginalization over the latent space. This is shown under Sec. A.4.2, and the resulting expectation exhibits a lack of dependence on sampling, as proven under Sec. A.4.3:
$\nabla_{\bm{f},\bm{\Phi}}\left[\log p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right]$ $=\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]+\nabla_{\bm{\Phi}}\left[\log p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right]\Bigg],$ (16)
2.4.1 Log-prior
$\log\Big(p_{\bm{f}}\left(\bm{z}\right)\Big)$ $=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}\log\left(\pi_{q,p}(z_{q,p})\right)=\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})+\log\left(\pi_{0}(z_{q,p})\right)-\log\left(Z_{q,p}\right)$ (17)
Tractable normalization is available through Eq. 6, however it may be more efficient to discard normalization in the learning gradient. In Sec. A.4.4, we derive the contrastive divergence learning gradient for KAEM, which faster to train and more typical of energy-based modeling:
$\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\left[\log p_{\bm{f}}\left(\bm{z}\right)\right]\Bigg]=\;$ $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]-\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg].$ (18)
Here, $\bm{z}$ is sampled from $p_{\bm{f}}(\bm{z})$ and $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$ and reflects a univariate element of the prior model. The expectation with respect to the posterior can be approximated using Importance Sampling (IS), outlined under Secs. 2.4.4 &  A.5.1, or a standard Monte Carlo estimator when using Langevin sampling methods. The prior expectation may always be approximated with a Monte Carlo estimator taken across samples:
$\mathbb{E}_{p_{\bm{f}}(\bm{z})}\Bigg[\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\Bigg]\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\nabla_{\bm{f}}\sum_{q=1}^{2n_{z}+1}\sum_{p=1}^{n_{z}}f_{q,p}(z_{q,p})\quad\text{where}\quad z_{q,p}^{(s)}\sim p_{q,p}(z).$ (19)
The gradient of Eq. 18 can therefore be viewed as a element-wise correction to the prior, incorporating information from the observations to align the prior with the posterior, thereby following the philosophy of empirical Bayes. The gradient can then be evaluated via autodifferentiation.
2.4.2 Log-likelihood
Depending on the task the log-likelihood can be modeled in multiple ways. For example, a Gaussian model might be adopted for images, and an autoregressive model for text. It can also be realized into a multimodal generative modeling framework using the principles outlined by yuan2024learningmultimodallatentgenerative.
$\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)$ $=\begin{cases}-\frac{\left\lVert\bm{x}^{(b)}-\bm{\tilde{x}}\right\rVert^{2}_{2}}{2\sigma_{\text{llhood}}^{2}}+\text{const},&\text{(static continuous data)}\\[5.0pt] \sum_{o=1}^{n_{x}}\sum_{c=1}^{K}x_{o,c}^{(b)}\log\tilde{x}_{o,c}+\text{const},&\text{(static discrete data)}\\[5.0pt] \sum_{t=1}^{T}\log\left(p_{\bm{\Phi}}(\bm{x}^{(b,t)}\mid\bm{x}^{(b,1)},\dots,\bm{x}^{(b,t-1)},\bm{z})\right),&\text{(sequential data)}\end{cases}$ (20)
Here, $\bm{\tilde{x}}$ is a generated sample from Eq. 9 corresponding to a single batch item, $\bm{x}^{(b)}$, $\sigma_{\text{llhood}}$ is a hyperparameter subject to assumptions regarding the variance of the data, and $K$ denotes the number of discrete classes. The learning gradient attributed to the log-likelihood can also be evaluated using autodifferentiation, after first approximating the posterior expectation with IS or MC estimators.
The code in Sec. A.1 also provides an option to augment the first likelihood model with an additional perceptual loss term (johnson2016perceptuallossesrealtimestyle) using features extracted from the pretrained VGG-16 CNN (simonyan2015deepconvolutionalnetworkslargescale). This could have yielded higher-fidelity images for Sec. 3.2, but it was not adopted in this study due to the computational overhead of Langevin sampling.
2.4.3 Prior inverse transform sampling
In Eqs. 9 &  19, $z_{q,p}$ is drawn from $p_{q,p}$ with inverse transform sampling, (ITS) (devroye1986nonuniform). ITS can be used to sample from a component of the latent density, $p_{q,p}$, made possible given the unique characteristics of KAEM. Specifically, the univariate, finite-grid structure of the KAN-based energy functions enables quadrature/interpolating integral approximations.
  1. 1.
    CDF inversion: Sampling occurs using the inverse cumulative density function (CDF): The CDF for a single component, $F_{\pi_{q,p}}$, is:
    $F_{\pi_{q,p}}(z^{\prime})$ $=\frac{\int_{z_{\text{min}}}^{z^{\prime}}\exp(f_{q,p}(z))\pi_{0}(z)\,dz}{Z_{q,p}},\quad Z_{q,p}=\int_{z_{\text{min}}}^{z_{\text{max}}}\exp(f_{q,p}(z))\pi_{0}(z)\,dz.$ (21)
    These two integrals can be conveniently and efficiently approximated using Gauss-Kronrod quadrature, (gauss-quad), given that $f_{q,p}$ is univariate and defined on a bounded grid:
    $G(z)$ $=\exp(f_{q,p}(z))\pi_{0}(z),\quad\int_{z_{\text{min}}}^{z^{\prime}}G(z)\,dz\approx\sum_{i=1}^{N_{\text{quad}}}w_{i}\,G(z_{i}^{\text{node}}),$ (22)
    where $z_{i}^{\text{node}}$ and $w_{i}$ are the Gauss-Legendre quadrature nodes and weights for the interval $z_{\text{min}},z^{\prime}]$. A random variable is uniformly drawn and shared across $q$, $u_{p}\sim\text{Unif}(0,1)$. The first interval where $F_{\pi_{q,p}}(z)$ is greater than $u_{p}$ scaled by $Z_{q,p}$ is identified:
    $k^{*}=\min\Bigg\{j\mid\sum_{i=1}^{j}w_{i}\,G(z_{i}^{\text{node}})\geq Z_{q,p}\cdot u_{p}\Bigg\}.$ (23)
  2. 2.
    Linear interpolation: The sample $z_{p}$ is interpolated within the subset, $z_{k^{*}-1}^{\text{node}},z_{k^{*}}^{\text{node}}]$:
    $z_{p}$ $\approx z_{k^{*}-1}^{\text{node}}+\left(z_{k^{*}}^{\text{node}}-z_{k^{*}-1}^{\text{node}}\right)\cdot\frac{u_{p}-F_{\pi_{q,p}}(z_{k^{*}-1}^{\text{node}})}{F_{\pi_{q,p}}(z_{k^{*}}^{\text{node}})-F_{\pi_{q,p}}(z_{k^{*}-1}^{\text{node}})}.$ (24)
The procedure ensures an efficient, exact and unbiased generation of samples, $z_{q,p}$, distributed according to $p_{q,p}$. It can be vectorized to generate samples from all prior elements.
2.4.4 Posterior Expectations with Importance Sampling
The log-marginal likelihood and its learning gradient presented in Eqs. 15 &  16, involve expectations with respect to the posterior $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$, which is potentially multimodal and complex. We propose Importance Sampling (IS) (goertzel1949quota) as an efficient method for estimating its expectation when datasets are simple or when task knowledge can be incorporated into the model to improve prior-posterior alignment and the quality of the IS estimator.
IS provides a quick approximation of the posterior expectation without requiring extensive exploration. In contrast, ULA, (introduced in Sec. 2.5.2) is more stochastic and exploratory, often yielding higher-quality samples from the posterior, when the prior is poorly matched.
The following procedure is elaborated in more detail under Sec. A.5.1:
  1. 1.
    Proposal distribution: $N_{s}$ independent samples, $\{\bm{z}^{(s)}\}_{s=1}^{N_{s}}$, are drawn from the proposal, $q(\bm{z})=p_{\bm{f}}(\bm{z})$, which the target, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})$ must be absolutely continuous with respect to. Sampling from the prior is tractable with the procedure outlined in Sec. 2.4.3.
  2. 2.
    Importance weights: The normalized importance weights are evaluated:
    $w_{\text{norm}}(\bm{z}^{(s)})=\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N_{s}}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\bm{\Phi}}(\bm{x}^{(b)}\mid\bm{z}^{(s)}))\Big).$ (25)
  3. 3.
    Effective Sample Size: To prevent weight degeneracy, resampling is conducted when the Effective Sample Size (ESS) falls below a threshold specified by a hyperparameter, $\gamma$:
    $ESS=\frac{1}{\sum^{N_{s}}_{s=1}w_{\text{norm}}(\bm{z}^{(s)})}<\gamma\,\cdot\,N_{s}.$ (26)
    Otherwise, the weights are taken without alteration, (i.e. the next step is skipped). When the weights are uniformly spread the ESS evaluates to $N_{s}$. High-variance weights with a low ESS indicate that many samples contribute minimally, resulting in inefficient sampling. Resampling focuses learning on samples that are more representative of the posterior distribution, which maximizes the usefulness of each parameter update.
  4. 4.
    Resample: If required, the population, $\{\bm{z}^{(s)},w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N_{s}}$, is resampled from. There are multiple ways to accomplish this, (douc2005comparisonresamplingschemesparticle). Systematic, stratified, and residual resampling are provided in the codebase under Sec. A.1, parallelized across samples. In this study, we adopted residual resampling, (explained under Sec. A.5.2).
  5. 5.
    Expectation estimation: After resampling, IS weights are uniformly distributed and the posterior expectation of an arbitrary function or density, $\rho(\bm{z})$, (e.g., the log-likelihood), is estimated with a Monte Carlo estimator:
    $\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)})}[\rho(\bm{z})]$ $\approx\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{\bm{z}^{(s)})}\approx\frac{1}{N_{s}}\sum_{s=1}^{N_{s}}\rho(\bm{z}^{(s)}_{\text{resampled}})$. (27)
This is used to estimate the log-marginal likelihood and its gradient in Eqs. 15 and 16. However, unlike ULA (discussed in Sec. 2.5.2 and demonstrated in Sec. 3.2), IS fails to support posterior exploration when the prior is poorly aligned, making it unsuitable for general-purpose tasks beyond those similar to the datasets presented in Secs. 3.1.
2.5 Population-based and Langevin Monte Carlo Posterior Sampling
To mitigate the poor convergence of multimodal ULA sampling, we introduce a training criterion that facilitates posterior annealing. Details on the cyclic annealing schedule adopted in this study, along with practical guidance on hyperparameter selection, are provided in Sec. A.3.
2.5.1 Thermodynamic Training Criterion
We inform our derivations using the Thermodynamic Integral. In place of the log-marginal likelihood outlined in Eq. 15, this uses the following representation derived in Sec. A.4.5:
$\log\left(p_{\bm{f},\bm{\Phi}}(\bm{x}^{(b)})\right)=\int_{0}^{1}\mathbb{E}_{p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)}\left[\log\left(p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)\right)\right]\,dt$ (28)
Here, $p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)$ has been introduced as the power posterior, which can be considered a representation of the original posterior, with a likelihood tempered by a temperature parameter, $t$. This is introduced by 2d186535-310b-3428-8c6f-e4ef298483d7 and organized for our purposes as follows:
$p_{\bm{f},\bm{\Phi}}(\bm{z}\mid\bm{x}^{(b)},t)$ $=\frac{p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\cdot p_{\bm{f}}(\bm{z})}{\bm{Z}_{t}},$
$\quad\text{where}\quad\bm{Z}_{t}=\mathbb{E}_{p_{\bm{f}}(\bm{z})}\left[p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\right]$ $=\int_{\mathcal{Z}}p_{\bm{\Phi}}\left(\bm{x}^{(b)}\mid\bm{z}\right)^{t}\,\,\mathbb{P}_{\bm{f}}(d\bm{z})$ (29)

### v9

@@ B.2 Kolmogorov-Arnold Network generator
Thus, a smooth transition between the latent and data space can be achieved by progressively doubling widths, eliminating the need for arbitrary choices while maintaining compliance with KART. Strict adherence to KART can be achieved for hidden layer $h_{l_{k},l_{k-1}}$ by applying either a sigmoid or affine transformation after layer $h_{l_{k-2},l_{k-3}}$ to shift its range back to $0,1]^{L_{k-1}}$. However, using sigmoid may lead to vanishing gradients and we suggest that function domains are less critical to KART than its structural bias and should not be prioritized over maintaining gradient flow and avoiding saturation.
Appendix C Importance sampling
C.1 Theory
Importance Sampling (IS) is a Monte Carlo that enables the estimation of expectations with respect to one probability measure using samples drawn from another. For a detailed exposition, see Kloek and van Dijk (1978).
The Radon-Nikodym derivative
IS involves rewriting the expectation of an arbitrary function or density, $\rho(\bm{z})$, under the target measure, $\mathbb{P}(\bm{z})$, in terms of a more tractable proposal measure, $\mathbb{Q}(\bm{z})$. Specifically, we use the Radon-Nikodym derivative $\frac{d\mathbb{P}}{d\mathbb{Q}}$ to rewrite the expectation of an arbitrary function of $\bm{z}$, denoted as $\rho(\bm{z})$:
$\mathbb{E}_{\mathbb{P}}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\,\mathbb{P}(d\bm{z})=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})\cdot\frac{d\mathbb{P}}{d\mathbb{Q}}\,\mathbb{Q}(d\bm{z}).$ (33)
In the context of KAEM, the target density is the posterior, $p_{f,\Phi}(\bm{z}\mid\bm{x})$, which is proportional to the product of the likelihood and the prior:
$p_{f,\Phi}(\bm{z}\mid\bm{x})\propto p_{\Phi}(\bm{x}\mid\bm{z})\cdot p_{f}(\bm{z}).$ (34)
The exact normalization constant, (the marginal likelihood $p_{f,\Phi}(\bm{x})$), is intractable, but IS circumvents the need to compute it explicitly. Instead, the expectation of $\rho(\bm{z})$ under the posterior can be expressed as:
$\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}[\rho(\bm{z})]=\int_{\mathcal{\bar{Z}}}\rho(\bm{z})w(\bm{z})\,q(\bm{z})d\bm{z},$ (35)
where the importance weights $w(\bm{z})$ are given by:
$w(\bm{z})=\frac{p_{f,\Phi}(\bm{z}\mid\bm{x})}{q(\bm{z})}=\frac{p_{\Phi}(\bm{x}\mid\bm{z})\cdot p_{f}(\bm{z})}{q(\bm{z})}.$ (36)
Proposal distribution and importance weights
A practical choice for the proposal density $q(\bm{z})$ is the prior, $p_{f}(\bm{z})$, given the availability of draws from the prior using the method outlined in Algorithm 1. This simplifies the importance weights, as the prior has a tractable form and covers the support of the posterior. This yields:
$w(\bm{z}^{(s)})=p_{\Phi}(\bm{x}\mid\bm{z}^{(s)})\propto\exp(\log(p_{\Phi}(\bm{x}\mid\bm{z}^{(s)}))).$ (37)
This reflects that the likelihood now directly informs how much weight to assign to each sample $\bm{z}$. Intuitively, samples that better explain the observed data $\bm{x}$, (as measured by the likelihood), are given higher importance. Using the definition of softmax, the normalized weights are given by:
$\frac{w(\bm{z}^{(s)})}{\sum_{r=1}^{N}w(\bm{z}^{(r)})}=\text{softmax}_{s}\Big(\log(p_{\Phi}(\bm{x}\mid\bm{z}^{(s)}))\Big).$ (38)
Resampling
The suitability of these weights depends on how well $q(\bm{z})$ matches the posterior. In KAEM, there is likely to be significant mismatch due to the complexity of the likelihood used in Eq. 8. Consequently, the weights will exhibit high variance, resulting in many samples contributing minimally. This reduces the Effective Sample Size (ESS) and introduces bias into the estimates.
To address this, the weights are resampled before proceeding with the posterior expectation. This is accomplished by using one of the methods outlined by Douc et al. (2005), whenever the ESS falls below a certain threshold determined by $\text{ESS}<\gamma*N$, where $\gamma$ is a hyperparameter controlling resample rate and $N$ is the number of samples. Resampling redistributes the sample weights, creating a new population with uniformly distributed weights while preserving the statistical properties of the original distribution. As training progresses and the prior becomes a better match for the posterior, the need for resampling decreases.
Importance Sampling estimator
To estimate the expectation, we draw $N$ independent samples $\{\bm{z}^{(s)}\}_{s=1}^{N}\sim q(\bm{z})=p_{f}(\bm{z})$ using the procedure outlined in Algorithm 1 and compute a weighted sum:
$\mathbb{E}_{p_{f,\Phi}(\bm{z}\mid\bm{x})}[\rho(\bm{z})]\approx\sum_{s=1}^{N}\rho(\bm{z}^{(s)})\cdot w_{\text{norm}}(\bm{z}^{(s)})\approx\frac{1}{N}\sum_{s=1}^{N}\rho(\bm{z}^{(s)}_{\text{resampled}})$. (39)
The estimator is unbiased, implying that its expectation matches the true posterior expectation. As $N\to\infty$, the approximation converges almost surely to the true value, provided that $q(\bm{z})$ sufficiently covers the posterior to ensure a well-conditioned Radon-Nikodym derivative. This estimator is used for the log-marginal likelihood and its gradient in Eqs. 4, since it provides a means of estimating expectations with respect to the posterior.
C.2 Residual resampling
In this study, we adopted residual resampling to filter the population of latent samples whenever the effective sample size (ESS) criterion in Algorith 2 is satisfied. Given a set of normalized weights $\{w_{\text{norm}}(\bm{z}^{(s)})\}_{s=1}^{N}$, the method deterministically replicates samples according to the integer part of $Nw_{\text{norm}}$, and uses multinomial sampling to allocate the remaining draws according to the residual mass.
Algorithm 3 Residual resampling (ESS-triggered)
0: Latrnt samples $\{z^{(s)}\}_{s=1}^{N}$ with normalized likelihood-based weights $\{w_{s}\}_{s=1}^{N}$ ($\sum_{s=1}^{N}w_{s}=1$); threshold fraction $\gamma\in(0,1]$.
0: Resampled $\{\tilde{z}^{(i)}\}_{i=1}^{N}$ and weights $\{\tilde{w}_{i}\}_{i=1}^{N}$.
1:ESS gating: $\mathrm{ESS}\leftarrow\Big(\sum_{s=1}^{N}w_{s}^{2}\Big)^{-1}$.
2:if $\mathrm{ESS}\geq\gamma N$ then
3:  for $i=1$ to NN do
4:   $\tilde{z}^{(i)}\leftarrow z^{(i)}$
5:   $\tilde{w}_{i}\leftarrow w_{i}$
6:  end for
7:  return $\{(\tilde{z}^{(i)},\tilde{w}_{i})\}_{i=1}^{N}$ {no resampling}
8:end if
9:Integer replication and residual mass:
10:for $s=1$ to NN do
11:  $r_{s}\leftarrow\lfloor Nw_{s}\rfloor$
12:  $m_{s}\leftarrow Nw_{s}-r_{s}$ {residual mass}
13:end for
[…]
15:Initialize index list: $\mathcal{I}\leftarrow[\,]$.
16:for $s=1$ to NN do
17:  Append $s$ to $\mathcal{I}$ exactly $r_{s}$ times
[…]
19:if $R>0$ then
20:  Normalize residual weights:
21:  for $s=1$ to NN do
22:   $\bar{w}_{s}\leftarrow m_{s}/R$
23:  end for
24:  Build CDF: $C_{0}\leftarrow 0$
25:  for $s=1$ to NN do
26:   $C_{s}\leftarrow C_{s-1}+\bar{w}_{s}$
27:  end for
28:  Draw remaining $R$ indices (multinomial):
29:  for $i=1$ to $R$ do
30:   Sample $u\sim\mathrm{Unif}(0,1)$
31:   Find smallest $k^{\star}$ such that $C_{k^{\star}}\geq u$
32:   Append $k^{\star}$ to $\mathcal{I}$
33:  end for
34:end if
35:Resample and reset weights:
36:for $i=1$ to NN do
37:  $\tilde{z}^{(i)}\leftarrow z^{(\mathcal{I}_{i})}$
38:  $\tilde{w}_{i}\leftarrow 1/N$
39:end for
40:return $\{(\tilde{z}^{(i)},\tilde{w}_{i})\}_{i=1}^{N}$.
Appendix D Guidance on annealing
D.1 Cyclic annealing